# Mitigating Perception Bias: A Training-Free Approach to Enhance LMM for Image Quality Assessment

### v1

@@ 4.2 Comparison with NR-IQA Models
Visualization. To verify the effectiveness of our method, we visualize our predicted results alongside the LMM predictions using the Q-Bench prompt on the KADID-10k dataset (first two rows) and the AIGC-3k dataset (second two rows). As shown in Fig. 4, we can observe that: 1) Despite comparable distortions and closely aligned MOS distributions in the first two rows, the LMM without our debiasing enhancement consistently assigns higher quality ratings to images in the first row over the second. This observation underscores the model’s high reliance on semantic content rather than low-level clues for quality assessment, revealing the presence of perceptual bias. 2) The bias varies by semantic content, affecting both natural and AI-generated images. In contrast, our debias strategy could effectively mitigate such bias, resulting in predictions that are more consistent with human ratings (i.e., MOSs).
4.3 Generalization on Other LMMs
In our method, we adopted the multimodal model mPLUG-Owl3 as our foundation model. To demonstrate the generalization capability of our enhancement strategy, we further validate it on another two LMMs inducing: BakLLaVA [26], and Qwen-VL [3]. As shown in Table 2, we could observe a consistent average performance gains can be achieved. Notably, the improvements observed on the LIVE Challenge and AGIQA-3k datasets suggest that semantic bias is widespread across diverse content types and such bias can be mitigated by our method efficiently. The promising generalization capability highlights the transformative potential of our bias-mitigation strategy and opens up exciting new avenues for developing training-free enhancement methods to fully harness the potential of LMMs for unseen tasks.
Table 4: Study of instructive prompt.
@@ 4.3 Generalization on Other LMMs
Q-Debias 0.794 0.790 0.838 0.863 0.717 0.753
Table 5: Study of different aggregation schemes.
@@
Q-Debias 0.794 0.790 0.838 0.863 0.717 0.753
4.4 Ablation Studies
Study of Conditional Images. As shown in Table 3, we compare our method with eight types of variants: (1) no conditional images, (2)-(5) each variant with a single image, distorted by one of the four distortion types individually, (6) two images with the top two performing distortion types, (7) three images with the top three performing distortion types, and (8) the conditional images with semantics that do not match the query image. To establish a representative baseline, we first select images from open-source datasets that do not share semantics with those in the LIVE Challenge, SIQAD, and AGIQA-3k datasets. Four specific distortions are then applied to these selected images to create an additional conditional image set. Finally, we randomly select four low-quality images from this set as the conditional images. The results reveal that: 1) Each distortion type contributes positively to our final image quality prediction, with the Saturate distortion achieving the highest performance on the natural images and generated images (LIVE Challenge and AGIQA-3k datasets). 2) Combining multiple conditional images leads to final performance improvements, highlighting the complementary roles of the four distortion types. 3) A significant performance drop can be witnessed when semantically inconsistent images are used as conditions, underscoring the critical importance of semantic correspondence in bias mitigation. This reveals that bias is highly semantic-specific; using conditional images with matching semantics to the query image enables remarkably accurate bias estimation and consistently enhances quality prediction.
Study of Instructive Prompt. In our method, the prompt serves an instructive role for the LMM, facilitating the propagation of bias mitigation from the conditional images to the query image. To verify its effectiveness, we compare our method against three prompt variants: (T1) Prompt Replacement: The entire prompt is replaced with “Rate the quality of the second image. Good or poor?” (T2) Bias Exposure Ablation: The phrase “The visual quality of the first image is poor” is removed from our prompt to examine the role of bias exposure, leading to the second prompt: “How about the visual quality of the second image? Good or poor?” (T3) Bias Mitigation Propagation Ablation: We delete the “How about” from the prompt to assess the impact of bias mitigation propagation to the query image, resulting in the third prompt: “The visual quality of the first image is poor. Rate the visual quality of the second image. Good or poor?” As shown in Table 4, the results reveal that: 1) Without our instructive prompt, even with the conditional images provided, bias cannot be effectively mitigated. 2) Without the explicit indication that the conditional images are of poor quality, the LMM fails to recognize the bias exposure, resulting in a marked performance drop. 3) The phrase “How about” suggests that the LMM should infer the query image’s quality based on the prior understanding that the conditional images are of poor quality. Without this phrase, the propagation of bias mitigation from the conditional images to the query image weakens noticeably. The best results are achieved when the full prompt is included, demonstrating the necessity of each instruction in our prompt.
@@ 4.4 Ablation Studies
Study of Aggregation Scheme. To aggregate the quality scores derived from different conditional images, we introduce a semantic similarity aggregation strategy. To assess its effectiveness, we compare our method with four alternative schemes: 1) Average Aggregation: Each of the four quality scores is assigned an equal weight during aggregation. 2) Quality Similarity Aggregation: We utilize the widely adopted FR-IQA model, LPIPS [63], to measure the quality similarity between the query image and each of its conditional images. These quality similarity scores are then treated as weights for the aggregation. 3) Winner-Takes-All: The final quality score is determined solely by the quality score obtained from the conditional image that exhibits the highest semantic similarity to the query image. The results are presented in Table 5, which reveal that: 1) The Average scheme results in a noticeable performance drop, suggesting that uniform weighting fails to account for the bias variations across different types of distortions. 2) The Quality Similarity scheme is also ineffective. The possible reason may lie that a higher quality similarity score does not always correspond to a higher semantic recognition for the LMM, due to perceptual discrepancy between the LMM and the human visual system. 3) The Winner-Takes-All scheme, though commonly used for score aggregation, demonstrates suboptimal performance as it fails to adequately capture the nuanced contributions of different conditional images. In comparison, our semantic similarity aggregation scheme delivers the best performance across datasets containing synthetic, authentic, and generative distortions, demonstrating its superior generalization on diverse image distortions.
5 Conclusion

### v2

@@ Introduction
Visualization. To verify the effectiveness of our method, we visualize our predicted results alongside the LMM predictions using the Q-Bench prompt on the KADID-10k dataset (first two rows) and the AIGC-3k dataset (second two rows).
As shown in Fig. 4, we can observe that: 1) Despite comparable distortions and closely aligned MOS distributions in the first two rows, the LMM without our debiasing enhancement consistently assigns higher quality ratings to images in the first row over the second. This observation underscores the model’s high reliance on semantic content rather than low-level clues for quality assessment, revealing the presence of perceptual bias. 2) The bias varies by semantic content, affecting both natural and AI-generated images. In contrast, our debias strategy could effectively mitigate such bias, resulting in predictions that are more consistent with human ratings (i.e., MOSs).
Generalization on other LMMs
In our method, we adopted the multimodal model mPLUG-Owl3 as our foundation model. To demonstrate the generalization capability of our enhancement strategy, we further validate it on another four LMMs inducing: BakLLaVA (Liu et al. 2024), Qwen-VL (Bai et al. 2023), LLaVA-OneVision (Li et al. 2024a), LLaVA-Interleave (Li et al. 2024b) and DeepSeek-VL2 (Wu et al. 2024). As shown in Table 2, we could observe a consistent average performance gains can be achieved. Notably, the improvements observed on the LIVE Challenge and AGIQA-3k datasets suggest that semantic bias is widespread across diverse content types and such bias can be mitigated by our method efficiently. The promising generalization capability highlights the transformative potential of our bias-mitigation strategy and opens up exciting new avenues for developing training-free enhancement methods to fully harness the potential of LMMs for unseen tasks.
Ablation Studies
Table 4: Ablation study of instructive prompt.
@@
Q-Debias 0.794 0.818 0.838 0.863 0.717 0.753
Study of Conditional Images. In our method, distortion types are carefully selected to degrade image quality while preserving semantic content. Specifically, we explore five types of distortions for conditional image construction: blur, noise, adverse weather conditions, brightness adjustment, and saturation modification. For blur distortion, we consider zoom blur, motion blur, and Gaussian blur, while for noise distortion, we examine Gaussian noise and spatter noise. Additionally, we synthesize significant snow, frost, and fog distortions, as humans can still recognize objects in images captured under adverse weather conditions. The study results are summarized in Table 3.
From Table 3, we observe that all distortion types enhance AI-generated images in the AGIQA-3k dataset, whereas only a subset (e.g., spatter noise and saturation modification) improves quality prediction for natural images in the LIVE Challenge dataset. This suggests that perception bias in AI-generated images is more pronounced, making performance gains more detectable. A potential explanation is that LMMs have been exposed to significantly fewer AI-generated images compared to natural images during training due to the vast historical disparity in dataset sizes. Moreover, different distortion types exhibit varying effectiveness in bias mitigation, underscoring the importance of careful distortion design, as bias levels are highly dependent on image semantics. To develop a generalized and effective bias mitigation strategy, we further explore potential combinations of the examined distortions. Given the exponential growth in possible combinations, we evaluate a four-distortion combination scheme, selecting the most effective distortion from each category. The results demonstrate that incorporating diverse distortion types yields the highest performance gains, highlighting their complementary roles in mitigating bias.
In our approach, distortions are applied directly to the query image to generate degraded versions while maintaining semantic consistency. To assess the necessity of semantic consistency, we further investigate the use of conditionally distorted images with mismatched semantics. Specifically, we construct an additional conditional image set by applying the four selected distortions to open-source images that do not share semantic content with those in the LIVE Challenge and AGIQA-3k datasets. For each quality inference, we randomly select four low-quality images from this set as conditional images. The results, presented in Exp. 14 of Table 3, show a significant performance drop when semantically inconsistent images are used, reinforcing the critical role of semantic alignment in bias mitigation. This finding highlights that bias is highly semantic-specific—leveraging conditionally degraded images that align with the query image semantics enables more accurate bias estimation and ultimately improves quality prediction.
It is important to note that our four-distortion combination represents just one potential bias mitigation strategy. Further exploration of different distortion combinations may yield superior performance. Nonetheless, our findings provide a strong foundation for refining and optimizing bias mitigation techniques in image quality assessment.
Study of Instructive Prompt. In our method, the prompt serves an instructive role for the LMM, facilitating the propagation of bias mitigation from the conditional images to the query image. To verify its effectiveness, we compare our method against three prompt variants: (T1) Prompt Replacement: The entire prompt is replaced with “Rate the quality of the second image. Good or poor?(T2) Bias Exposure Ablation: The phrase “The visual quality of the first image is poor” is removed from our prompt to examine the role of bias exposure, leading to the second prompt: “How about the visual quality of the second image? Good or poor?(T3) Bias Mitigation Propagation Ablation: We delete the “How about” from the prompt to assess the impact of bias mitigation propagation to the query image, resulting in the third prompt: “The visual quality of the first image is poor. Rate the visual quality of the second image. Good or poor?” As shown in Table 4, the results reveal that: 1) Without our instructive prompt, even with the conditional images provided, bias cannot be effectively mitigated. 2) Without the explicit indication that the conditional images are of poor quality, the LMM fails to recognize the bias exposure, resulting in a marked performance drop. 3) The phrase “How about” suggests that the LMM should infer the query image’s quality based on the prior understanding that the conditional images are of poor quality. Without this phrase, the propagation of bias mitigation from the conditional images to the query image weakens noticeably. The best results are achieved when the full prompt is included, demonstrating the necessity of each instruction in our prompt.
Table 5: Ablation study of different aggregation schemes.
@@
Q-Debias 0.794 0.818 0.838 0.863 0.717 0.753
@@
Study of Aggregation Scheme. To aggregate the quality scores derived from different conditional images, we introduce a semantic similarity aggregation strategy. To assess its effectiveness, we compare our method with four alternative schemes: 1) Average Aggregation: Each of the four quality scores is assigned an equal weight during aggregation. 2) Quality Similarity Aggregation: We utilize the widely adopted FR-IQA model, LPIPS (Zhang et al. 2018), to measure the quality similarity between the query image and each of its conditional images. These quality similarity scores are then treated as weights for the aggregation. 3) Winner-Takes-All: The final quality score is determined solely by the quality score obtained from the conditional image that exhibits the highest semantic similarity to the query image. The results are presented in Table 5, which reveal that: 1) The Average scheme results in a noticeable performance drop, suggesting that uniform weighting fails to account for the bias variations across different types of distortions. 2) The Quality Similarity scheme is also ineffective. The possible reason may lie that a higher quality similarity score does not always correspond to a higher semantic recognition for the LMM, due to perceptual discrepancy between the LMM and the human visual system. 3) The Winner-Takes-All scheme, though commonly used for score aggregation, demonstrates suboptimal performance as it fails to adequately capture the nuanced contributions of different conditional images. In comparison, our semantic similarity aggregation scheme delivers the best performance across, demonstrating superior generalization on diverse image distortions.
Conclusion
@@
Acknowledgments
The work was supported by the National Natural Science Foundation of China under Grant No. 62401214.